Erosion costs. Fat Tire Bicycle Company currently sells 40,000 bicycles per year. The current bike is a standard​ balloon-tire bike selling for ​$120​, with a production and shipping cost of ​$25. The company is thinking of introducing an​ off-road bike with a projected selling price of ​$410 and a production and shipping cost of ​$225. The projected annual sales for the​ off-road bike are 17,000. The company will lose sales in​ fat-tire bikes of 7500 units per year if it introduces the new​ bike, however. What is the erosion cost from the new​ bike? Should Fat Tire start producing the​ off-road bike?
Answer #1

Erosion cost from the new bike should be equal to=7500*(120-25)
=$712,500
Incremental earnings from off road bikes=17000*(410-225)=$3,145,000

As incremental earnings is more than erosion cost, Fat Tire should start producing the off-road bike